OpenStudy (anonymous):

A hemispherical tank has a total surface area of 243pieft2 this icclude the circular base area . Find the volume

OpenStudy (tkhunny):

\(Volume\;=\;\dfrac{4}{3}\pi r^{3}\) \(Surface\;Area\;=\;4\pi r^{2}\) All that really is left is to figure out if the tank has a lid that we should be counting.